OpenStudy (anonymous):

The students at Midtown Middle school sold flowers as a fundraiser in September and October. In October, they charged $1.50 for each flower. The October price was a 20% increase of the September price. Part A What was the price of the flowers in September?

OpenStudy (anonymous):

@Study312 I got 1.20

OpenStudy (study312):

Funny I just did this. I am in Flvs 7th grade math too

OpenStudy (anonymous):

Lol what teacher? also please don't give me the answer directly

OpenStudy (study312):

Mr. Niebauer, who do you have?

OpenStudy (anonymous):

Mrs.Robinson

OpenStudy (study312):

I have had Mr.Niebauer for 2 years

OpenStudy (anonymous):

I got that September earned more than October by 30$

OpenStudy (anonymous):

Wow. I used to have Mr. Reese but I think he got promoted

OpenStudy (study312):

0.2 • 1.50 = 0.3 0.3 + 1.50 = 1.80

OpenStudy (study312):

Can you give me part 2 of the question?

OpenStudy (anonymous):

But it says that October was the increase so wouldn't September cost less?

OpenStudy (anonymous):

Part B The seventh-grade class earned 40% of the selling price of each flower. In September, they sold 900 flowers. In October, they sold 700 flowers. The seventh-grade class earned more money in September/october by 30/70/190

OpenStudy (study312):

so 0.4 • 1.80 = 0.72 0.72 - 1.80 = 1.08

OpenStudy (anonymous):

Omg I did it wrong again! Idk what I was thinking but I First multiplied by the number of flowers THEN I subtracted the 40%

OpenStudy (study312):

1.08 • 900 = 972 1.50 • 0.4 = 0.6 1.50 - 0.6 = 0.9 0.9 • 700 = 630

OpenStudy (study312):

its subtracted by 40% first because it says "The seventh-grade class earned 40% of the selling price of EACH flower."

OpenStudy (study312):

972 - 630 = 342

OpenStudy (study312):

actually I think I am wrong, I don't see 342 as an option

OpenStudy (study312):

Can you show how you did it?

OpenStudy (anonymous):

Ohhh. See what I was doing was doing it in the order I thought it should be in, also I think that you have to subtract 20% from 1.50 instead of adding 20%

OpenStudy (anonymous):

Okay

OpenStudy (study312):

it said it increased by 20%

OpenStudy (anonymous):

No it said the October flowers increased from the september

OpenStudy (study312):

so lets try subtracting 20% and see what we get

OpenStudy (anonymous):

Okay let me try it

OpenStudy (study312):

ok

OpenStudy (study312):

My teacher told me it was 1.80 so I got the 2nd part wrong

OpenStudy (study312):

or maybe 1.80 is wrong

OpenStudy (anonymous):

I got 1.20 the first time so 1.20 * 0.40 =0.48 1.50 * 0.40 =0.60 0.48*900= 432 0.60 * 700=420 432-420= 12

OpenStudy (anonymous):

Nope not right either :/

OpenStudy (anonymous):

@LegendarySadist

OpenStudy (study312):

BRB

OpenStudy (anonymous):

Okay

OpenStudy (anonymous):

Do you see what we are doing wrong?

OpenStudy (study312):

back

OpenStudy (study312):

idk

OpenStudy (anonymous):

Hi I think that Legendary is seeing what we did wrong

OpenStudy (anonymous):

The very first step is wrong. If there is a 20% increase that means it is 120% compared to the first one. So \[\large \sf O=1.2S \huge \rightarrow \large 1.5=1.2S \huge \rightarrow \large S=\frac{1.5}{1.2}\]

OpenStudy (anonymous):

So 1.5 divided by 1.2?

OpenStudy (study312):

= 12.5

OpenStudy (anonymous):

=1.25

OpenStudy (anonymous):

Ohhhhhh

OpenStudy (anonymous):

yeah, 1.5 divided by 1.2 which is 1.25

OpenStudy (study312):

That makes more sense

OpenStudy (anonymous):

Yeah thank you :D Let me try to get the right answer...

OpenStudy (anonymous):

September= 450 October= 420

OpenStudy (study312):

450 - 420 = 30

OpenStudy (anonymous):

So September earns more by 30!

OpenStudy (study312):

Thx Legendary.

OpenStudy (anonymous):

Yes thank you so much both of you :D

OpenStudy (anonymous):

Ye no problem

OpenStudy (anonymous):

Study I will PM you if it was right

OpenStudy (study312):

ok